# ReplaceMe: Network Simplification via Depth Pruning and Transformer Block Linearization

### v3

@@ Abstract
We introduce ReplaceMe, a generalized training-free depth pruning method that effectively replaces transformer blocks with a linear operation, while maintaining high performance for low compression ratios. In contrast to conventional pruning approaches that require additional training or fine-tuning, our approach requires only a small calibration dataset that is used to estimate a linear transformation, which approximates the pruned blocks. The estimated linear mapping can be seamlessly merged with the remaining transformer blocks, eliminating the need for any additional network parameters. Our experiments show that ReplaceMe consistently outperforms other training-free approaches and remains highly competitive with state-of-the-art pruning methods that involve extensive retraining/fine-tuning and architectural modifications. Applied to several large language models (LLMs), ReplaceMe achieves up to 25% pruning while retaining approximately 90% of the original model’s performance on open benchmarks—without any training or healing steps, resulting in minimal computational overhead (see Fig. 1). We provide an open-source library implementing ReplaceMe alongside several state-of-the-art depth pruning techniques, available at https://github.com/mts-ai/ReplaceMe.
[…]
Figure 1: Comparison of the proposed LLM compression method with state-of-the-art UIDL. Subplots illustrate (a) compression time, (b) environmental impact ($\text{CO}_{2}$ emissions and energy consumption), and (c) performance accuracy relative to the uncompressed baseline. Our approach attains the shortest compression time, lowest energy use, and reduced emissions, while achieving the highest accuracy, demonstrating superior efficiency, sustainability, and effectiveness over existing methods.
@@ 2.1 Layers selection
$i^{*}=\operatorname*{arg\,min}_{i}D\left(\text{L}_{i},\text{L}_{i+n}\right).$ (4)
We evaluated various distance metrics $D(\cdot)$ and found that the cosine distance is particularly effective in identifying nearly optimal layers for pruning. This observation aligns with findings reported in recent studies [12]. In the supplementary material A.10, we present results of an exhaustive brute-force layer selection, which confirm that cosine distance consistently identifies optimal or near-optimal layers for removal. Additionally, we provide a comparative analysis of the $L_{2}$ distance metric to further validate our choice.
@@ 2.2 Linear Transform Estimation
To compensate for the pruned transformer blocks, we leverage a small set of calibration data to compute activations directly before and after the removal point. Using these activations, we estimate an appropriate linear transformation that accurately approximates the final output of the pruned blocks. Depending on the selected criterion, this estimation can be performed using either analytical or numerical approaches, enabling precise modeling of the omitted layers.
@@
A classical solution for Eq. 5 arises when setting $h(.)$ as the $L_{2}$-distance and is obtained by solving a least-squares (LS) problem:
@@
As discussed in Section 2.1, our ablation study comparing various distance functions to assess the importance of transformer blocks, revealed that the cosine distance is the most effective in identifying the least significant blocks. Motivated by these results, we have further used the cosine distance as the objective function to estimate the optimal linear transformation. In this case, the optimization problem takes the form:
$\mathbf{T}^{*}=$ $\operatorname*{arg\,min}_{\mathbf{T}}\mathrm{cosine\_distance}\left(\mathbf{M}_{i}\cdot\mathbf{T}+\mathbf{Y}_{i},\mathbf{L}_{i+n}\right)$
@@
where we use the notation $\mathbf{M}_{i,k}\in\mathbb{R}^{d}$ to denote the $k$-th row of a matrix $\mathbf{M}_{i}\in\mathbb{R}^{N\times d}$, which we then represent as a column vector. Here, the cosine distance is calculated per token $k$ and aggregated over all $N$ tokens. Unlike the $L_{2}$-distance formulation, this objective does not admit a closed-form solution, and thus a numerical optimization approach is required. In our experiments, we have utilized the Adam [18] optimization algorithm. Furthermore, an ablation study involving various alternative numerical solvers can be found in the supplementary material A.6.
@@
$\mathbf{T}^{*}=\operatorname*{arg\,min}_{\mathbf{T}}\mathrm{cosine\_distance}\left(\mathbf{M}_{i}\cdot\mathbf{T},\mathbf{L}_{i+n}-\mathbf{Y}_{i}\right).$ (8)
@@
Merging the Linear Transformation
Once the optimal transformation $\mathbf{T}^{*}$ has been estimated, our approach allows it to be incorporated into the MLP layer of the $i$-th transformer block, where it is merged with the weight matrix of the second FFN layer. Consequently, the overall architecture of the model remains unchanged, except for the removal of the “non-effective” transformer blocks.
@@ 2.3 Regularization
$\mathbf{T}^{*}=$ $\operatorname*{arg\,min}_{\mathbf{T}}h(\mathbf{M}_{i}\cdot\mathbf{T}+\mathbf{Y}_{i};\mathbf{L}_{i+n})+\alpha\cdot R(\mathbf{T}),$ (9)
@@ 2.4 Multiple Linear Transforms
The proposed ReplaceMe method can be easily extended to be applied on multiple non-overlapping groups of blocks within the model, estimating a separate linear transformation for each group (multi-LT). This approach provides flexibility in achieving the desired performance metrics, even under high compression ratios. Furthermore, if some of the selected groups of blocks are consecutive, they can be merged into a single block with one corresponding linear transformation. We refer to this method as non-consecutive Multi-linear transformations (Multi_LT_nc), experimental analyses further validating Multi_LT_nc are detailed in the Appendix A.16.
@@ 3.1 Experimental setup
In our experiments, we have focused primarily on LLaMA-2-7B and LLaMA-3-8B-Instruct models, while also reporting results for Qwen2.5-7B and Falcon-11B for a comparative analysis. Results for additional models are provided in the supplementary material A.7. For numerical estimation of the linear transform, we used Adam optimizer with LR $1e^{-4}$ and batch size 1024, iterating for 10 epochs over the calibration data. In Table 1 we present results on different benchmarks that have been widely used in previous research. These benchmarks have been introduced in the following works: CMNLI [50], HellaSwag [51], PIQA [2], CHID [53], WSC [22], MMLU [15], CMMLU [23], Race-High/Middle [20], C3 [41]. Additionally, we benchmarked  ReplaceMe using well-established public datasets, namely Winogrande [37], BoolQ [4], OpenBookQA [29], SciQ [49], and Lambada OpenAI [32]. For all benchmarks except Lambada OpenAI, we report accuracy as the evaluation metric, along with the average accuracy across all benchmarks. For Lambada OpenAI, we report perplexity.
@@ 3.2 Comparison with other structured-pruning methods
Table 1: Comparing other pruning methods after healing and our training free approach ReplaceMe, * indicates that the numbers are taken from streamline paper [3]. After compressing Llama 2 7B with 25% compression ratio. signifies that the model was trained following pruning, whereas indicates that the model is training-free.
@@
Llama 3 8B Instruct [7] 3.11 0.7 100
SVD-LLM [47] None 29.90 0.59 85.3
LLMPruner [27] None 12.31 0.60 85.3
UIDL [12] Identity 2216.96 0.58 82.5
ReplaceMe(ours) Linear (LS) 20.23 0.63 89.9
ReplaceMe(ours) Linear (Cosine) 15.88 0.63 90.9
ReplaceMe(ours) Multi_LT_nc (Cosine) 13.95 0.63 90.0
@@
2222444466660.550.550.550.550.60.60.60.60.650.650.650.650.70.70.70.7AccuracyLlama 3 8B instructUIDLReplaceMe Adam00555510101010151515150.550.550.550.550.60.60.60.60.650.650.650.650.70.70.70.7Falcon 2 11B2222444466660.550.550.550.550.60.60.60.60.650.650.650.650.70.70.70.7Qwen 2.5 7B instruct
@@
Finally, we note that at high compression ratios, applying a healing process becomes necessary, as linear transformations alone are insufficient to fully recover performance. Although ReplaceMe continues to outperform UIDL under these conditions, a healing phase is required at extreme compression levels to maintain model effectiveness.
@@ 3.3.1 Ablation on calibration data
Furthermore, in the supplementary material (Section A.12), we present a comprehensive analysis of the impact of calibration dataset size, including the minimum number of samples required to achieve stable model performance. We also introduce a masking-based augmentation technique designed to maintain robust performance even when only a limited subset of data samples is available, ensuring computational efficiency under resource-constrained conditions.
@@ 4 Related Work
Model pruning [21, 13] has been at the frontier of deep-learning research since the early developments of this field. It has found practical applications not only in model size reduction but also in enhancing the interpretability of the models under study. The same holds true for pruning LLMs.
@@ A.2 Closed-form solution for $L_{2}$ Distance
$0\displaystyle 2\mathbf{M}_{i}^{\top}\mathbf{M}_{i}\cdot\mathbf{T}+2\mathbf{M}_{i}^{\top}(\mathbf{Y}_{i}-\mathbf{L}_{i+n})=$ 0
@@ A.8 Computational Efficiency Analysis
As demonstrated in Figure 4, a comparative analysis of our two proposed methodologies is presented. The LS approach offers significant advantages in terms of cost-efficiency, speed, and the reduced demand for computational resources and memory. However, this method incurs a slight decline in performance metrics. Thus, the choice between these methods should be informed by the specific requirements and available hardware resources of the user.
@@ A.11 Cosine Distance Approximation
$\mathbf{T}^{*}=\operatorname*{arg\,min}_{\mathbf{T}}\mathrm{cosine\_distance}\left(\mathbf{M}_{i}\cdot\mathbf{T}+\mathbf{Y}_{i},\mathbf{L}_{i+n}\right),$ (14)
@@
$\mathbf{T}^{*}=\operatorname*{arg\,min}_{\mathbf{T}}\mathrm{cosine\_distance}\left(\mathbf{M}_{i}\cdot\mathbf{T},\mathbf{L}_{i+n}-\mathbf{Y}_{i}\right).$ (15)
Table 10: Comparison between accurate cosine distance and approximated version. cosine loss ($m,\,l$$y$) denotes that we subtracted the attention output of $i_{th}$ block from the activation of the$(i+n)_{th}$ block.
@@
Llama 3 8B instruct cosine loss ($m,\,l$$y$) 8 slim_orca no training 15.88 0.634
Llama 3 8B instruct cosine loss ($m+y,\,l$) 8 slim_orca no training 16.63 0.630
Qwen 2.5 7B instruct cosine loss ($m,\,l$$y$) 7 slim_orca no training 7.92 0.591
Qwen 2.5 7B instruct cosine loss ($m+y,\,l$) 7 slim_orca no training 10.60 0.580
Llama 3 8B instruct Multi_A cosine loss ($m,\,l$$y$) 8 slim_orca no training 13.95 0.628
Llama 3 8B instruct Multi_A cosine loss ($m+y,\,l$) 8 slim_orca no training 13.12 0.630
@@ A.12 Data ablation
0.10.10.10.$\cdot 10^{4}$000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Calibration SamplesAverage AccuracyAverage Accuracy vs. Calibration SamplesLSCosineLlama3 8B (Baseline)0.10.10.10.$\cdot 10^{4}$1010101020202020Calibration SamplesPerplexityPerplexity vs. Calibration Samples
@@
The linear transformation matrix has shape $d\times d$, requiring approximately $N=d^{2}$ tokens for reliable estimation. With a per-sample sequence length of $S$, this translates to at least $d^{2}/S$ samples. For example, in LLaMA 3 8B, where $d=4096$ and $S=1024$, about 16,000 samples are theoretically needed. However, as seen in Figure 6, accuracy remains stable even with as few as 1,000 samples, suggesting robustness to sample size. That said, perplexity continues to improve with more data, indicating better model confidence and predictive quality.

### v4

# Dmitriy Shopkhoev

@@ Abstract
We introduce ReplaceMe, a generalized training-free depth pruning method that effectively replaces transformer blocks with a linear operation, while maintaining high performance for low compression ratios. In contrast to conventional pruning approaches that require additional training or fine-tuning, our approach requires only a small calibration dataset that is used to estimate a linear transformation, which approximates the pruned blocks. The estimated linear mapping can be seamlessly merged with the remaining transformer blocks, eliminating the need for any additional network parameters. Our experiments show that ReplaceMe consistently outperforms other training-free approaches and remains highly competitive with state-of-the-art pruning methods that involve extensive retraining/fine-tuning and architectural modifications. Applied to several large language models (LLMs), ReplaceMe achieves up to 25% pruning while retaining approximately 90% of the original model’s performance on open benchmarks—without any training or healing steps, resulting in minimal computational overhead. We provide an open-source library implementing ReplaceMe alongside several state-of-the-art depth pruning techniques, available at https://github.com/mts-ai/ReplaceMe.
@@ 2.1 Layers selection
$i^{*}=\operatorname*{arg\,min}_{i}h\left(\text{L}_{i},\text{L}_{i+n}\right).$ (4)
We evaluated various distance metrics $h(\cdot)$ using activations computed on a small calibration dataset, with the impact of the dataset explored in Section 3.3.1. For each candidate cut, the distance between hidden states before and after the cut is calculated and then averaged across the calibration dataset to obtain a robust layer-importance estimate. We found that the cosine distance is particularly effective in identifying nearly optimal layers for pruning. This observation aligns with findings reported in recent studies [13]. In the supplementary material A.10, we present results of an exhaustive brute-force layer selection, which confirm that cosine distance consistently identifies optimal or near-optimal layers for removal. Additionally, we provide a comparative analysis of the $L_{2}$ distance metric to further validate our choice.
@@ 2.2 Linear Transform Estimation
To compensate for the pruned transformer blocks, we leverage the same set of calibration data to compute activations directly before and after the removal point. Using these activations, we estimate an appropriate linear transformation that accurately approximates the final output of the pruned blocks. Depending on the selected criterion, this estimation can be performed using either analytical or numerical approaches, enabling precise modeling of the omitted layers.
@@
A classical solution for Eq. 5 arises when setting $h(\cdot)$ as the $L_{2}$-distance and is obtained by solving a least-squares (LS) problem:
@@
As discussed in Section 2.1, our ablation study on various distance functions to assess the importance of transformer blocks, revealed that the cosine distance is the most effective in identifying the least significant blocks. Motivated by these results, we have further used the cosine distance as the objective function to estimate the optimal linear transformation. In this case, the optimization problem takes the form:
$\mathbf{T}^{*}=$ $\operatorname*{arg\,min}_{\mathbf{T}}\mathrm{cos}\left(\mathbf{M}_{i}\cdot\mathbf{T}+\mathbf{Y}_{i},\mathbf{L}_{i+n}\right)$
@@
where $\mathrm{cos(\cdot,\cdot)}$ denotes the cosine distance between the two input vectors. We use the notation $\mathbf{M}_{i,k}\in\mathbb{R}^{d}$ to denote the $k$-th row of a matrix $\mathbf{M}_{i}\in\mathbb{R}^{N\times d}$, which we then represent as a column vector. Here, the cosine distance is calculated per token $k$ and aggregated over all $N$ tokens. Unlike the $L_{2}$-distance formulation, this objective does not admit a closed-form solution, and thus a numerical optimization approach is required. In our experiments, we have utilized the Adam [19] optimization algorithm. Furthermore, an ablation study involving various alternative numerical solvers can be found in the supplementary material A.6.
@@
$\mathbf{T}^{*}=\operatorname*{arg\,min}_{\mathbf{T}}\mathrm{cos}\left(\mathbf{M}_{i}\cdot\mathbf{T},\mathbf{L}_{i+n}-\mathbf{Y}_{i}\right).$ (8)
@@
Fusing the Linear Transformation
Once the optimal transformation $\mathbf{T}^{*}$ has been estimated, our approach allows it to be incorporated into the MLP layer of the $i$-th transformer block. Since the learned transformation and the pre-cut MLP block represent a sequential linear operators, they can be algebraically composed into a single equivalent linear transformation, allowing $\mathbf{T}^{*}$ to be fused with the weight matrix of the down projection linear layer of the MLP block. Consequently, the overall architecture of the model remains unchanged, except for the removal of the “non-effective” transformer blocks.
@@ 2.3 Regularization
$\mathbf{T}^{*}=\$$\operatorname*{arg\,min}_{\mathbf{T}}h(\mathbf{M}_{i}\cdot\mathbf{T}+\mathbf{Y}_{i};\mathbf{L}_{i+n})+\alpha\cdot R(\mathbf{T}),$ (9)
@@ 2.4 Multiple Linear Transforms
The proposed ReplaceMe method can be easily extended to be applied on multiple non-overlapping groups of blocks within the model, estimating a separate linear transformation for each group (multi-LT). This approach provides flexibility in achieving the desired performance metrics, even under high compression ratios. Furthermore, if some of the selected groups of blocks are consecutive, they can be merged into a single block with one corresponding linear transformation. We refer to this method as non-consecutive Multi-linear transformations (Multi_LT_NC). Experimental analyses further validating Multi_LT_NC are detailed in the Appendix A.16.
@@ 3.1 Experimental setup
In our experiments, we have focused primarily on LLaMA-2-7B and LLaMA-3-8B-Instruct models, while also reporting results for Qwen2.5-7B and Falcon-11B for a comparative analysis. Results for additional models are provided in the supplementary material A.7. For numerical estimation of the linear transform, we used Adam optimizer with LR $1e^{-4}$ and batch size 1024, iterating for 10 epochs over the calibration data. In Table 1 we present results on different benchmarks that have been widely used in previous research. These benchmarks have been introduced in the following works: CMNLI [56], HellaSwag [57], PIQA [2], CHID [59], WSC [24], MMLU [16], CMMLU [25], Race-High/Middle [22], C3 [46]. Additionally, we benchmarked  ReplaceMe using well-established public datasets, namely Winogrande [41], BoolQ [4], OpenBookQA [31], SciQ [55], and Lambada OpenAI [35]. For all benchmarks except Lambada OpenAI, we report accuracy as the evaluation metric, along with the average accuracy across all benchmarks. For Lambada OpenAI, we report perplexity.To calculate the environmental impact we used CodeCarbon [6], a well-known tool to estimate $CO_{2}$ emissions and power consumption
@@ 3.2 Comparison with other structured-pruning methods
EmissionsTime0505010010015015011113.53.56.66.6107107169169Rel. to LSEnvironmental Normalized ComparisonOurs (LS)  Ours (Cosine)  UIDLOurs (Cosine)Ours (LS)UIDL85859090959510010092.592.589.989.990.390.3Relative AccuracyRelative Accuracy
Figure 2: Comparison of the proposed LLM compression method with state-of-the-art UIDL. Subplots illustrate (a) compression time and environmental impact ($\text{CO}_{2}$ emissions), and (b) performance accuracy relative to the uncompressed baseline. Our approach attains the shortest compression time and reduced emissions, while achieving the highest accuracy, demonstrating superior efficiency, sustainability, and effectiveness over existing methods.
@@
Table 1: Comparing other pruning methods after healing and our training free approach ReplaceMe, * indicates that the numbers are taken from streamline paper [3]. After compressing Llama 2 7B with 25% compression ratio. signifies that the model was trained following pruning, whereas indicates that the model is training-free. We also report the relative performance to the original, unpruned model (RP).
@@
Llama 3 8B Instruct [8] 3.11 0.7 100%
SVD-LLM [53] None 29.90 0.59 85.3%
LLMPruner [29] None 12.31 0.60 85.3%
UIDL [13] Identity 2216.96 0.58 82.5%
ReplaceMe(ours) Linear (LS) 20.23 0.63 89.9%
ReplaceMe(ours) Linear (Cosine) 15.88 0.63 90.9%
ReplaceMe(ours) Multi_LT_NC (Cosine) 13.95 0.63 90.0%
@@
2244660.550.550.60.60.650.650.70.7AccuracyLlama 3 8B instructUIDLReplaceMe Adam055101015150.550.550.60.60.650.650.70.7Falcon 2 11B2244660.550.550.60.60.650.650.70.7Qwen 2.5 7B instruct
@@
Finally, we note that at high compression ratios, applying a healing process becomes necessary, as linear transformations alone are insufficient to fully recover performance. Details are provided in Table 3. Although ReplaceMe continues to outperform UIDL under these conditions, a healing phase is required at extreme compression levels to maintain model effectiveness.
Compression ratio 12.5% 25% 37.5% 50%
UIDL 59.9 50.4 41 30.4
Ours 64.6 60.7 51 38.6
Table 3: Average accuracy for our method and UIDL at higher compression ratios is compared across the Race, Winogrande, PIQA, BoolQ, OpenBookQA, and SciQ benchmarks.
@@ 3.3.1 Ablation on calibration data
Furthermore, in the supplementary material (Section A.12), we present a comprehensive analysis of the impact of calibration dataset size, including the minimum number of samples required to achieve stable model performance. We also introduce a masking-based augmentation technique designed to maintain robust performance even when only a limited subset of data samples is available, ensuring computational efficiency under resource-constrained conditions. In addition, Section A.17 presents results with task-specific calibration datasets, showing that using data from the same task as the evaluation target (e.g., SciQ) further improves accuracy compared to general-purpose calibration data such as SlimOrca.
@@ 4 Related Work
Model pruning [23, 14] has been at the frontier of deep-learning research since the early developments of this field. It has found practical applications not only in model size reduction but also in enhancing the interpretability of the models under study. Earlier studies on transformer-based language models, such as BERT, demonstrated that these models are highly compressible [42, 40]. The same observation holds for pruning LLMs.
@@ A.2 Closed-form solution for $L_{2}$ Distance
$2\mathbf{M}_{i}^{\top}\mathbf{M}_{i}\cdot\mathbf{T}+2\mathbf{M}_{i}^{\top}(\mathbf{Y}_{i}-\mathbf{L}_{i+n})=0$
@@ A.8 Computational Efficiency Analysis
As demonstrated in Fig. 4, a comparative analysis of our two proposed methodologies is presented. The LS approach offers significant advantages in terms of cost-efficiency, speed, and the reduced demand for computational resources and memory. However, this method incurs a slight decline in performance metrics. Thus, the choice between these methods should be informed by the specific requirements and available hardware resources of the user.
@@ A.10 Block Selection Analysis
In Fig. 6, we present the detailed results of sequential layer removal and compare the cosine distance with other similarity measures: CKA [20] and CCA [32]. We observe that the highest accuracy corresponds to the lowest cosine distance, which confirms the validity of using cosine distance as a block selection criterion.
It is also important to note that the optimal pruning range depends on both the architecture and the number of layers removed. Previous works have also observed that layers form natural groups [49]. We conducted a similar experiment and plotted the cosine distance across all layers in Fig. 7.
Figure 6: Comparison of cosine distance, CCA, and CKA across all 8-layer pruning configurations in Llama3 8B.
Figure 7: Average cosine distance between hidden states of layers for Llama 3 8B. Each cell $(i,j)$ shows the mean cosine distance between the output vectors of $layer_{i}$ and $layer_{j}$, computed across 256 samples of SlimOrca.
@@ A.11 Cosine Distance Approximation
$\mathbf{T}^{*}=\operatorname*{arg\,min}_{\mathbf{T}}\mathrm{cos}\left(\mathbf{M}_{i}\cdot\mathbf{T}+\mathbf{Y}_{i},\mathbf{L}_{i+n}\right),$ (14)
@@
$\mathbf{T}^{*}=\operatorname*{arg\,min}_{\mathbf{T}}\mathrm{cos}\left(\mathbf{M}_{i}\cdot\mathbf{T},\mathbf{L}_{i+n}-\mathbf{Y}_{i}\right).$ (15)
@@
Llama 3 8B instruct approx. loss 8 slim_orca no training 15.88 0.634
Llama 3 8B instruct orig. loss 8 slim_orca no training 16.63 0.630
Qwen 2.5 7B instruct approx. loss 7 slim_orca no training 7.92 0.591
Qwen 2.5 7B instruct orig. loss 7 slim_orca no training 10.60 0.580
Llama 3 8B instruct Multi_A approx. loss 8 slim_orca no training 13.95 0.628
Llama 3 8B instruct Multi_A orig. loss 8 slim_orca no training 13.12 0.630
Table 11: Comparison between the original loss and the approximated version. The approximated version makes use of the cosine distance between the transformed MLP output and the residual of the activation of $(i+n)$-th block and the attention of the $i$-th block.
@@ A.12 Data ablation
[…]
@@
The linear transformation matrix has shape $d\times d$, requiring approximately $N=d^{2}$ tokens for reliable estimation. With a per-sample sequence length of $S$, this translates to at least $d^{2}/S$ samples. For example, in LLaMA 3 8B, where $d=4096$ and $S=1024$, about 16,000 samples are theoretically needed. However, as seen in Fig 8, accuracy remains stable even with as few as 1,000 samples, suggesting robustness to sample size. That said, perplexity continues to improve with more data, indicating better model confidence and predictive quality.